About once every 30 minutes, a geyser known as Old Faceful projects water 12.0 m straight up into the air. Use g = 9.80 m/s^2, and take atmospheric pressure to be 101.3 kPa. The density of water is 1000 kg/m^3.
What is the speed of the water when it emerges from the ground?
_______ m/s

About once every 30 minutes, a geyser known as Old Faceful projects water 12.0 m straight up into the air. Use g = 9.80 m/s^2, and take atmospheric pressure to be 101.3 kPa. The density of water is 1000 kg/m^3.
Assuming the water travels to the surface through a narrow crack that extends 6.00 m below the surface, and that the water comes from a chamber with a large cross-sectional area, what is the pressure in the chamber?
_______ kPa
Answer #1

Answon- a) By applying gauge pressione P2-P1 = & Pv² = egh tev1000 x9.8x12 12 - 117600X2 1000 V= 15.34 mis - P2-P1 = 2 ev teg